Grade 12th passMechanics

One proton beam enters a magnetic field of 10-4 m/s normally, sp. charge = 1011 C/kg, velocity = 109 m/s. What is the radius of the circle describe by it ?

Abin Santhosh
8 Years agoGrade 12th pass

2 Answers

Arun
8 Years ago
Dear student
 
I didn’t get your numerical details. Please clear all the details.
I will be happy to help you.
though you can use 
mv^2 / r = q*v*B

Regards
Arun (askIITians forum expert)
sana
7 Years ago
One proton beam enters a magnetic field of 10-4 m/s normally, sp. charge = 1011 C/kg, velocity = 109 m/s. What is the radius of the circle describe by it ?